the position of a particle moving along an x axis is given by x= 15t^2-5t^3, where x is in meters and t is in seconds. Determine a) the position ,b) the velocity,and c) the acceleration of the particle at t= 7.00s d) what is the maximum positive coordinate reached by the particle
Answer #1

2 Now, Velocity (V) V30 -1s t2 ond acziaraton (av di. 301-1 t) -30-30t 30 01 t) 악 2. = 735-1715 980 m = 210-チ35 V -525 m acreす 30t-152 au unan pahce wll each at manimum distane v01 0 30-15t-o 3) 15 2 -5(2)--5(2) = = 15x4-5X8 2260-90 (L = 20 m man